# The KSBA moduli space of stable log Calabi–Yau surfaces

### 23 Feb 2024 v1

@@ 7.1.1. Central fibers of polarized mirror families
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. We denote by $0_{\mathcal{X}}$ the unique torus fixed point of the affine toric variety $S_{\mathcal{X}}=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$.
@@
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. The central fiber $(Y_{0_{\mathcal{X}}},L_{0,\mathcal{X}})$ of the polarized mirror family $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ over $0_{\mathcal{X}}\in S_{\mathcal{X}}$ is the union toric surfaces with intersection complex $(P,\mathscr{P})$. In particular, the irreducible components of $Y_{0_{\mathcal{X}}}$ are copies of $\mathbb{P}^{2}$ indexed by the integral points $p\in P_{\mathbb{Z}}$, and $L_{0,\mathcal{X}}$ restricted to each irreducible component is $\mathcal{O}_{\mathbb{P}^{2}}(1)$.
@@
This follows from [44, §2.1], where central fibers of mirror families given by the intrinsic mirror construction are described explicitly. ∎
@@
As shown in §6.3, the wall-crossing transformations induced by the walls of the canonical scattering diagram are trivial modulo $J$, except for the contributions of the kinks corresponding to the finitely double curves $C$ of $\mathcal{X}_{0}$ with $F\cdot C=0$. The results follows by [40, §2.1] – see also [38, §6.2.1] and [16, §2]. ∎
@@ 7.1.2. Proof of finite generation
For every face $H$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$, the affine toric variety $\mathrm{Spec}\,\mathbb{C}[H]$ is a toric stratum of the affine toric variety $\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$. We denote by $T_{H}:=\mathrm{Spec}\,\mathbb{C}[H^{\mathrm{gp}}]$ the corresponding open dense torus in $\mathrm{Spec}\,\mathbb{C}[H]$. The second half of the proof of the following Lemma is inspired by [40, Proposition 6.6] (see also [63, Lemma 16.6] and [70, Proposition 4.19]).
@@
Let $G$ be the face of $H$ spanned by the finitely many curves $C$ such that $F\cdot C=0$, and denote by $J:=H\setminus G$ the corresponding monoid ideal. By Lemma 7.2, $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]/J$ is a Stanley-Reisner algebra over $\mathbb{C}[H]/J\simeq\mathbb{C}[G]$, and so in particular a finitely generated $\mathbb{C}[H]/J$-algebra. In particular, there exists a finite set $\{\vartheta_{p}\}_{p\in I}$ of theta functions generating $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]/J$ as $\mathbb{C}[H]/J$-algebra. We will show that $\{\vartheta_{p}\}_{p\in I}$ actually generate $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$ as $\mathbb{C}[H]$-algebra. Denoting by $(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$ the $\mathbb{C}[H]$-subalgebra of $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$ generated by $\{\vartheta_{p}\}_{p\in I}$, we have to show that $(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}=\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$.
@@
Fix $w,h\in\mathbb{Z}_{\geq 0}$. We prove that every monomial $t^{\beta}\vartheta_{p}$ in $V_{v,w}$ is actually in $(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$ by decreasing induction on $-F\cdot\beta$. First note that, as $(hP)_{\mathbb{Z}}$ is finite, there are finitely many possible values for $-\varphi_{F}(p)$, and so, as $w=-\varphi_{F}(p)-F\cdot\beta$ is fixed, there are finitely many possible values for $-F\cdot\beta$ for $t^{\beta}\vartheta_{p}$ in $V_{v,w}$. Hence, the result is vacuously true for $-F\cdot\beta$ large enough. For the induction step, as $\{\vartheta_{q}\}_{q\in C(P)_{\mathbb{Z}}}$ generates $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]/J$, one can write $\vartheta_{p}=a+b$, with $a\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$ and $b\in J(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])$, and with both $a$ and $b$ homogeneous with respect to the $\mathbb{C}^{\star}$-action and of height $h$. Hence, we have $t^{\beta}\vartheta_{p}=t^{\beta}a+t^{\beta}b$, with $t^{\beta}a\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$, and $t^{\beta}b\in V_{w,h}$. Actually, as $b\in J(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])$, for every monomial $t^{\beta^{\prime}}\vartheta_{q}$ in $b$, we have $-F\cdot\beta^{\prime}\geq 1$. Hence, every monomial in $t^{\beta}b$ is of the form $t^{\beta+\beta^{\prime}}\vartheta_{q}$ with $-F\cdot(\beta+\beta^{\prime})>-F\cdot\beta$. It follows from the induction hypothesis that $t^{\beta}b\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$, and so finally that $t^{\beta}\vartheta_{p}\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$. ∎
@@ 7.2. Generic KSBA stability of the mirror family
In this section, we prove in Theorem 7.7 the KSBA stability of the restriction of the extended mirror family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}_{\mathcal{X}}^{\mathrm{sec}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{sec}})\rightarrow S_{\mathcal{X}}^{\mathrm{sec}}$ to the dense torus in $S_{\mathcal{X}}^{\mathrm{sec}}$.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. Then, the mirror family $(\mathcal{Y}_{\mathcal{X}},\mathcal{D}_{\mathcal{X}})\longrightarrow S_{\mathcal{X}}$ is a projective flat family of semi-log-canonical surfaces $(Y_{t},D_{t})$ such that $K_{Y_{t}}+D_{t}=~{}0$.
@@
First note that $(\mathcal{Y}_{\mathcal{X}},\mathcal{D}_{\mathcal{X}})\longrightarrow S_{\mathcal{X}}$ is flat since the algebra $\mathcal{R}_{\mathcal{X}}$ of theta functions is a free $NE(\mathcal{X}/\overline{\mathcal{X}})$-module. On the other hand, by Theorem 7.4, the algebra $\mathcal{R}_{\mathcal{X}}$ is finitely generated, and so the morphism $\mathcal{Y}_{\mathcal{X}}=\mathrm{Proj}\,\mathcal{R}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ is projective.
To show that the fiber $(Y_{t},D_{t})$ over $t\in S_{\mathcal{X}}$ is semi-log-canonical and $K_{Y_{t}}+D_{t}=0$, consider the smallest toric stratum $\mathrm{Spec}\,\mathbb{C}[H]$ of $S_{\mathcal{X}}$ containing $t$, corresponding to a face $H$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$. In particular, $t$ is contained in the dense torus orbit $T_{H}:=\mathrm{Spec}\,\mathbb{C}[H^{\mathrm{gp}}]$ in $\mathrm{Spec}\,\mathbb{C}[H]$. As in the proof of Lemma 7.3, there is a corresponding birational contraction $f_{H}:\mathcal{X}\rightarrow\mathcal{X}_{H}$ such that $H=NE(\mathcal{X}/\mathcal{X}_{H})$, and the restriction $\mathcal{Y}_{\mathcal{X}}|_{T_{H}}$ is independent of the choice of the crepant resolution $\mathcal{X}\rightarrow\mathcal{X}_{H}$. Hence, by Lemma 6.12, one can assume that there exists a good divisor $F$ on $\mathcal{X}$ over $\mathcal{X}_{H}$. Let $G$ be the face of $H$ spanned by the curves $C$ with $F\cdot C=0$, and denote $J:=H\setminus G$. Consider the one-parameter subgroup $\mathbb{C}^{\star}\subset T_{0}$ associated to $F$ as in §6.5. Consider the corresponding orbit $\mathbb{C}^{\star}\in T_{H}$ passing through $t$. Its closure in $\mathrm{Spec}\mathbb{C}[H]$ is an affine line $\mathbb{A}^{1}$ with $0\in\mathrm{Spec}\,\mathbb{C}[G]\simeq\mathrm{Spec}\,\mathbb{C}[H]/J$. By Lemma 6.27, the restriction of $(\mathcal{Y}_{\mathcal{X}},\mathcal{D}_{\mathcal{X}})$ to this line is a $\mathbb{C}^{\star}$-equivariant family. Hence, by the proof of [42, Lemma 8.33], it is enough to show Theorem 7.5 for the central fiber of this family over $0\in\mathrm{Spec}\mathbb{C}[G]\simeq\mathrm{Spec}\mathbb{C}[H]/J$. By Lemma 7.2, this central fiber is a fiber of a Mumford toric degeneration, and then the result follows from [6]. ∎
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. Then, the extended mirror family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}})\longrightarrow S^{\mathrm{sec}}_{\mathcal{X}}$ is a projective flat family of semi-log-canonical surfaces $(Y_{t},D_{t})$ such that $K_{Y_{t}}+D_{t}=0$.
@@
For the fibers of $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}})\longrightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$, the result follows from Theorem 7.5 applied to the various projective crepant resolutions of $\overline{\mathcal{X}}$. By Theorem 6.28, any fiber of $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}})\longrightarrow S^{\mathrm{sec}}_{\mathcal{X}}$ is related by $T_{0}$-action, and so is isomorphic, to a fiber of $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}})\longrightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$, and so the general result follows. ∎
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. For every $t\in S^{\mathrm{sec}}_{\mathcal{X}}$, let $(Y_{t},D_{t}+\epsilon C_{t})$ be the fiber over $t$ of the extended mirror family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\longrightarrow S^{\mathrm{sec}}_{\mathcal{X}}$. Then, the following holds:
@@
    For every $t\in S^{\mathrm{sec}}_{\mathcal{X}}$, if the surface $Y_{t}$ is normal, then the pair $(Y_{t},D_{t})$ is log canonical, and the pair $(Y_{t},D_{t}+\epsilon\,C_{t})$ is KSBA stable.
@@
    If $t$ is contained in the dense torus orbit $T_{\mathcal{X}}:=\mathrm{Spec}\,\mathbb{C}[N_{1}(\mathcal{X}/\overline{\mathcal{X}})]\subset S^{\mathrm{sec}}_{\mathcal{X}}$, then the surface $Y_{t}$ is normal, the pair $(Y_{t},D_{t})$ is log canonical, and the pair $(Y_{t},D_{t}+\epsilon\,C_{t})$ is KSBA stable.
@@
We first prove i). By assumption, the surface $Y_{t}$ is normal. By Theorem 7.6, the pair $(Y_{t},D_{t})$ is semi-log-canonical, and so log canonical since $Y_{t}$ is normal. Hence, by [49, Proposition 7.2], to prove that $(Y_{t},D_{t}+\epsilon C_{t})$ is semi-log-canonical, it is enough to show that $C_{t}$ does not contain any 0-dimensional stratum of $D_{t}$. By Remark 6.29, the divisor $C_{t}$ is defined an equation of the form $\sum_{p\in P_{\mathbb{Z}}\setminus I}\vartheta_{p}=0$ for some subset $I\subset P_{\mathbb{Z}}$ consisting of integral points corresponding to either 1-surfaces or 2-surfaces in $\mathcal{X}_{0}$. By Equations (6.8)-(6.9), we have $\vartheta_{p}|_{D_{t}}=0$ if $p\notin\partial P_{\mathbb{Z}}$. On the other hand, by Lemma 6.24, $D_{t}$ is a cycle of rational curves, with 00-dimensional strata $x_{v}$ in one-to-one correspondence with the integral points $v$ of $\partial P$ corresponding to the 0-surfaces in $\mathcal{X}_{0}$. Moreover, the theta functions $\vartheta_{p}$ with $p\in\partial P$ restricts to toric monomials on the irreducible components of $D_{t}$. It follows that for every 00-dimensional stratum $x_{v}$ of $D_{t}$, we have $\vartheta_{p}(x_{v})\neq 0$ if $p=v$, and $\vartheta_{p}(x_{v})=0$ if $p\in P_{\mathbb{Z}}\setminus\{v\}$. Hence, we have $\sum_{p\in P_{v}}\vartheta_{p}(x_{v})\neq 0$, and so $x_{v}\notin C_{t}$. This concludes the proof of i).

### 30 Mar 2024 v2

@@ 7.1.1. Central fibers of polarized mirror families
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. The central fiber $(Y_{0_{\mathcal{X}}},L_{0,\mathcal{X}})$ of the polarized mirror family $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ over the unique torus fixed point $0_{\mathcal{X}}$ of the affine toric variety $S_{\mathcal{X}}=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$ is the union of toric surfaces with intersection complex $(P,\mathscr{P})$. In particular, the set of irreducible components of $Y_{0_{\mathcal{X}}}$ is in one-to-one correspondence with $P_{\mathbb{Z}}$, all the irreducible components of $Y_{0_{\mathcal{X}}}$ are isomorphic to $\mathbb{P}^{2}$, and the line bundle $L_{0,\mathcal{X}}$ restricted to each irreducible component of $Y_{0_{\mathcal{X}}}$ is isomorphic to $\mathcal{O}_{\mathbb{P}^{2}}(1)$.
@@
This follows from [45, §2.1], where the central fibers of mirror families given by the intrinsic mirror construction are described explicitly. ∎
@@
As shown in §6.3, the wall-crossing transformations induced by the walls of the canonical scattering diagram are trivial modulo $J$, except for the contributions of the kinks corresponding to the finitely double curves $C$ of $\mathcal{X}_{0}$ with $F\cdot C=0$. The results follows by [41, §2.1] – see also [39, §6.2.1] and [16, §2]. ∎
@@ 7.1.2. Proof of finite generation
For every face $H$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$, the affine toric variety $\mathrm{Spec}\,\mathbb{C}[H]$ is a toric stratum of the affine toric variety $\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$. We denote by $T_{H}:=\mathrm{Spec}\,\mathbb{C}[H^{\mathrm{gp}}]$ the corresponding open dense torus in $\mathrm{Spec}\,\mathbb{C}[H]$.
@@
Let $G$ be the face of $H$ spanned by the finitely many curves $C$ such that $F\cdot C=0$, and denote by $J:=H\setminus G$ the corresponding monoid ideal. By Lemma 7.2, $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]/J$ is a Stanley-Reisner algebra over $\mathbb{C}[H]/J\simeq\mathbb{C}[G]$, and so in particular a finitely generated $\mathbb{C}[H]/J$-algebra. In particular, there exists a finite set $\{\vartheta_{p}\}_{p\in I}$ of theta functions generating $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]/J$ as $\mathbb{C}[H]/J$-algebra. We will show that $\{\vartheta_{p}\}_{p\in I}$ actually generate $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$ as $\mathbb{C}[H]$-algebra. Denoting by $(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$ the $\mathbb{C}[H]$-subalgebra of $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$ generated by $\{\vartheta_{p}\}_{p\in I}$, we have to show that $(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}=\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$. To do this, we follow the path of the proof of [41, Proposition 6.6] (see also [64, Lemma 16.6] and [71, Proposition 4.19]).
@@
Fix $w,h\in\mathbb{Z}_{\geq 0}$. We prove that every monomial $t^{\beta}\vartheta_{p}$ in $V_{v,w}$ is actually in $(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$ by decreasing induction on $-F\cdot\beta$. First note that, as $(hP)_{\mathbb{Z}}$ is finite, there are finitely many possible values for $-\varphi_{F}(p)$, and so, as $w=-\varphi_{F}(p)-F\cdot\beta$ is fixed, there are finitely many possible values for $-F\cdot\beta$ for $t^{\beta}\vartheta_{p}$ in $V_{v,w}$. Hence, the result is immediate for large enough values of $-F\cdot\beta$. For the induction step, as $\{\vartheta_{q}\}_{q\in C(P)_{\mathbb{Z}}}$ generates $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]/J$, one can write $\vartheta_{p}=a+b$, with $a\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$ and $b\in J(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])$, and with both $a$ and $b$ homogeneous with respect to the $\mathbb{C}^{\star}$-action and of height $h$. Hence, we have $t^{\beta}\vartheta_{p}=t^{\beta}a+t^{\beta}b$, with $t^{\beta}a\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$, and $t^{\beta}b\in V_{w,h}$. Actually, as $b\in J(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])$, for every monomial $t^{\beta^{\prime}}\vartheta_{q}$ in $b$, we have $-F\cdot\beta^{\prime}\geq 1$. Hence, every monomial in $t^{\beta}b$ is of the form $t^{\beta+\beta^{\prime}}\vartheta_{q}$ with $-F\cdot(\beta+\beta^{\prime})>-F\cdot\beta$. It follows from the induction hypothesis that $t^{\beta}b\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$, and so finally that $t^{\beta}\vartheta_{p}\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$. ∎
@@ 7.2. Generic KSBA stability of the mirror family
In this section, we prove in Theorem 7.7 the KSBA stability of the restriction of the extended mirror family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}_{\mathcal{X}}^{\mathrm{sec}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{sec}})\rightarrow\mathcal{S}_{\mathcal{X}}^{\mathrm{sec}}$ to the dense torus in $\mathcal{S}_{\mathcal{X}}^{\mathrm{sec}}$.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. Then, the mirror family $(\mathcal{Y}_{\mathcal{X}},\mathcal{D}_{\mathcal{X}})\rightarrow S_{\mathcal{X}}$ is a projective flat family of semi-log-canonical surfaces $(Y_{t},D_{t})$ such that $K_{Y_{t}}+D_{t}=~{}0$.
@@
First note that $(\mathcal{Y}_{\mathcal{X}},\mathcal{D}_{\mathcal{X}})\rightarrow S_{\mathcal{X}}$ is flat since the algebra $\mathcal{R}_{\mathcal{X}}$ of theta functions is a free $NE(\mathcal{X}/\overline{\mathcal{X}})$-module. On the other hand, by Theorem 7.4, the algebra $\mathcal{R}_{\mathcal{X}}$ is finitely generated, and so the morphism $\mathcal{Y}_{\mathcal{X}}=\mathrm{Proj}\,\mathcal{R}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ is projective.
To show that the fiber $(Y_{t},D_{t})$ over $t\in S_{\mathcal{X}}$ is semi-log-canonical and $K_{Y_{t}}+D_{t}=0$, consider the smallest toric stratum $\mathrm{Spec}\,\mathbb{C}[H]$ of $S_{\mathcal{X}}$ containing $t$, corresponding to a face $H$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$. In particular, $t$ is contained in the dense torus orbit $T_{H}:=\mathrm{Spec}\,\mathbb{C}[H^{\mathrm{gp}}]$ in $\mathrm{Spec}\,\mathbb{C}[H]$. As in the proof of Lemma 7.3, there is a corresponding birational contraction $f_{H}:\mathcal{X}\rightarrow\mathcal{X}_{H}$ such that $H=NE(\mathcal{X}/\mathcal{X}_{H})$, and the restriction $\mathcal{Y}_{\mathcal{X}}|_{T_{H}}$ is independent of the choice of the crepant resolution $\mathcal{X}\rightarrow\mathcal{X}_{H}$. Hence, by Lemma 6.12, we can assume that there exists a good divisor $F$ on $\mathcal{X}$ over $\mathcal{X}_{H}$. Let $G$ be the face of $H$ spanned by the curves $C$ with $F\cdot C=0$, and denote $J:=H\setminus G$. Consider the one-parameter subgroup $\mathbb{C}^{\star}\subset T_{0}$ associated to $F$ as in §6.5. Consider the corresponding orbit $\mathbb{C}^{\star}\in T_{H}$ passing through $t$. Its closure in $\mathrm{Spec}\,\mathbb{C}[H]$ is an affine line $\mathbb{A}^{1}$ with $0\in\mathrm{Spec}\,\mathbb{C}[G]\simeq\mathrm{Spec}\,\mathbb{C}[H]/J$. By Lemma 6.27, the restriction of $(\mathcal{Y}_{\mathcal{X}},\mathcal{D}_{\mathcal{X}})$ to this line is a $\mathbb{C}^{\star}$-equivariant family. Hence, by the proof of [43, Lemma 8.33], it is enough to show Theorem 7.5 for the central fiber of this family over $0\in\mathrm{Spec}\,\mathbb{C}[G]\simeq\mathrm{Spec}\,\mathbb{C}[H]/J$. By Lemma 7.2, this central fiber is a fiber of a Mumford toric degeneration. Hence, the result follows from [6]. ∎
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. Then, the extended mirror family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ is a projective flat family of semi-log-canonical surfaces $(Y_{t},D_{t})$ such that $K_{Y_{t}}+D_{t}=0$.
@@
For the fibers of $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}})\rightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$, the result follows from Theorem 7.5 applied to the various projective crepant resolutions of $\overline{\mathcal{X}}$. By Theorem 6.28, any fiber of $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ is related by $T_{0}$-action, and so is isomorphic, to a fiber of $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}})\rightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$, and so the general result follows. ∎
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. For every $t\in\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$, let $(Y_{t},D_{t}+\epsilon\,C_{t})$ be the fiber over $t$ of the extended mirror family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$. Then, the following holds:
@@
    For every $t\in\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$, if the surface $Y_{t}$ is normal, then the pair $(Y_{t},D_{t})$ is log canonical, and the pair $(Y_{t},D_{t}+\epsilon\,C_{t})$ is KSBA stable.
@@
    If $t$ is contained in the dense torus orbit $T_{\mathcal{X}}:=\mathrm{Spec}\,\mathbb{C}[N_{1}(\mathcal{X}/\overline{\mathcal{X}})]\subset\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$, then the surface $Y_{t}$ is normal, the pair $(Y_{t},D_{t})$ is log canonical, and the pair $(Y_{t},D_{t}+\epsilon\,C_{t})$ is KSBA stable.
@@
We first prove i). By Theorem 7.6, the pair $(Y_{t},D_{t})$ is semi-log-canonical, and so log canonical since the surface $Y_{t}$ is assumed to be normal. Hence, by [50, Proposition 7.2], to prove that $(Y_{t},D_{t}+\epsilon\,C_{t})$ is semi-log-canonical, it is enough to show that $C_{t}$ does not contain any 0-dimensional stratum of $D_{t}$. By Remark 6.29, the divisor $C_{t}$ is defined by an equation of the form $\sum_{p\in P_{\mathbb{Z}}\setminus I}\vartheta_{p}=0$ for some subset $I\subset P_{\mathbb{Z}}$ consisting of integral points corresponding to either 1-surfaces or 2-surfaces in $\mathcal{X}_{0}$. By Equations (6.8)-(6.9), we have $\vartheta_{p}|_{D_{t}}=0$ if $p\notin\partial P_{\mathbb{Z}}$. On the other hand, by Lemma 6.24, $D_{t}$ is a cycle of rational curves, with 00-dimensional strata $x_{v}$ in one-to-one correspondence with the integral points $v$ of $\partial P$ corresponding to the 0-surfaces in $\mathcal{X}_{0}$. Moreover, the theta functions $\vartheta_{p}$ with $p\in\partial P$ restrict to toric monomials on the irreducible components of $D_{t}$. It follows that for every 00-dimensional stratum $x_{v}$ of $D_{t}$, we have $\vartheta_{p}(x_{v})\neq 0$ if $p=v$, and $\vartheta_{p}(x_{v})=0$ if $p\in P_{\mathbb{Z}}\setminus\{v\}$. Hence, we have $\sum_{p\in P_{v}}\vartheta_{p}(x_{v})\neq 0$, and so $x_{v}\notin C_{t}$. This concludes the proof of i).